5 years ago Sue was 5 times as old as her sister Sue is three times as old as her sister is now how old is Sue
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

5 years ago Sue was 5 times as old as her sister Sue is three times as old as her sister is now how old is Sue
----
S = Sue's age
R = sister's age
---
S-5 = 5*(R-5)
S = 3R
----
Sub for S in the 1st eqn
3R - 5 = 5*(R-5)
Solve for R
then solve for S

Answer by stanbon(75887)

5 years ago Sue was 5 times as old as her sister Sue is three times as old as her sister is now how old is Sue
------
Equations:
S-5 = 5(s-5)
S = 3s
---------
Substitute for "s" and solve for "S"::
S-5 = 5(S/3 - 5)
S-5 = (5/3)S - 25
(2/3)S = 20
S = 30 yrs (Sue's age now)
------------
Cheers,
Stan H.
-----------